# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ Abstract
Quantum systems, fabricated across various spatial scales from nano to micrometers, are very delicate and naturally sensitive to the variations of their environment. These features make them excellent candidates for serving as sensors with wide range of applications. The size versatility as well as their sensitivity, allow quantum sensors to achieve unprecedented precision with ultra-high spatial resolution. The main power of quantum sensors is achieved when the probe is composed of several particles. In this situation, quantum features such as entanglement contribute in enhancing the precision of quantum sensors beyond the capacity of classical sensors, a phenomenon that is referred to as quantum enhanced sensitivity. Originally, quantum sensing was formulated for non-interacting particles which are prepared in a special form of maximally entangled states. These probes are extremely sensitive to decoherence and any interaction between particles is detrimental to their performance. An alternative framework for quantum sensing has been developed exploiting quantum many-body systems, where the interaction between particles plays a crucial role. In this review, we investigate different aspects of the latter approach for quantum metrology and sensing. Many-body probes have been used for sensing purposes in both equilibrium and non-equilibrium scenarios. Quantum criticality has been identified as a resource for achieving quantum enhanced sensitivity in both of these scenarios. In equilibrium, various types of criticalities, such as first order, second order, topological, and localization phase transitions have been exploited for sensing purposes. In non-equilibrium scenarios, quantum enhanced sensitivity has been discovered for Floquet, dissipative, and time crystal phase transitions. While each type of these criticalities, either in equilibrium or non-equilibrium scenarios, has its own characteristics, the presence of one feature is crucial for achieving quantum enhanced sensitivity and that is energy/quasi-energy gap closing. In non-equilibrium quantum sensing, time becomes another parameter which can affect the sensitivity of the probe. Typically, the sensitivity enhances as the probe evolves in time. Unlike criticality-based quantum sensors, the criteria which allows non-equilibrium probes to achieve quantum enhanced sensitivity is not fully explored. In this review, we provide an overview on recent progresses on different aspects of quantum metrology and sensing with many-body systems.
@@ I Introduction
The rapidly emerging quantum technologies are expected to revolutionize our lives in the coming decades Dowling and Milburn (2003). In a broad sense, quantum technologies are developed in three directions, namely quantum simulation and computation Lloyd (1996); Georgescu et al. (2014), quantum communication Zeilinger (1999); Gisin and Thew (2007), and quantum sensing Degen et al. (2017). Thanks to advancements in fabricating ultra-precise quantum devices, the progress in all these three aspects have significantly been accelerated. Quantum sensing Degen et al. (2017) is one of the pillars of quantum technologies which allow for developing a new generation of sensors for detecting gravitational, magnetic and electric fields as well as estimating physical quantities with unprecedented precision, well beyond the capability of conventional classical sensors. The application prospect of quantum sensors is very wide and includes mining Crawford et al. (2021), environmental monitoring  Lanzagorta et al. (2015), biological imaging Wu et al. (2016a); Taylor and Bowen (2016); Aslam et al. (2023), space exploration Kaltenbaek et al. (2021), radar technology Maccone and Ren (2020); Karsa et al. (2024); Casariego et al. (2023), searching for fundamental particles Brady et al. (2022); Backes et al. (2021); Kang et al. (2024); Wang et al. (2022a); Jiang et al. (2021a) — the list goes on. There are several reasons that make quantum sensors superior over their classical counterparts. First, the even smaller sizes of quantum sensors, which can be fabricated at the atomic scales, allows for excellent spatial resolution Schirhagl et al. (2014); Zhang et al. (2021). Second, the natural delicacy of quantum superposition allows quantum sensors to detect very weak signals Hong et al. (2024); Jiang et al. (2021b). Third, a combination of quantum particles can benefit from a unique property of quantum physics, namely entanglement, which allows for quantum enhanced precision in comparison with a classical sensor using the same amount of resources Giovannetti et al. (2004, 2011, 2006).
@@
In order to quantify the precision of a sensor for estimating an unknown parameter $\theta$, one needs to specify a proper figure of merit. In case of maximal-likelihood estimation theory, the imprecision of the estimation, as quantified by the variance of an unbiased max-likelihood estimator, is bounded through Cramér-Rao inequality Cramer (1946); Rao (1992) by $1/MI(\theta)$, where $M$ is the number of samples and $I(\theta)$ is a quantity called Fisher information, which will be defined and discussed later. Dependence on $M$ is a direct consequence of central limit theorem, therefore the most important quantity which bounds the precision of a sensor is the Fisher information. In fact, in estimation theory, Fisher information is the conventional figure of merit for evaluating the performance of a sensor. Every sensing procedure exploits some resources such as time, particle number, system size, etc. The resource efficiency of a given sensor is determined by the scaling of Fisher information with respect to those resources. In other words, such scaling shows how the precision is enhanced by using more resources. In the absence of quantum features, the Fisher information at best scales linearly with the resource, which is known as standard (or shot-noise) limit. However, exploiting quantum features such as superposition and entanglement may result in super-linear scaling of the Fisher information with respect to the resource. This is known as quantum enhanced sensitivity and is one of the key properties that makes quantum sensors superior over their classical counterparts.
@@
The notion of quantum enhanced sensitivity was first discovered in an interferometric setup introduced in the seminal paper of Giovannetti, Lloyd and Maccone Giovannetti et al. (2004). In this paper, the authors show that a special form of entangled states, such as Greenberger–Horne–Zeilinger (GHZ) states, can enhance the precision of detecting a phase shift quadratically with respect to the probe size. Optical setups have since been used to experimentally confirm this effect quite extensively (Mitchell et al., 2004; Pezzé et al., 2007; Ono et al., 2013). Even for matter-based platforms, the effect was tested experimentally in ion-trap systems Leibfried et al. (2004) and then extended to various physical platforms like superconducting qubits Wang et al. (2019) and nitrogen vacancy centers Bonato et al. (2016). The original formulation Giovannetti et al. (2004) relies on the GHZ-type of entangled states which are difficult to create and extremely sensitive to decoherence and particle loss. In addition, in this strategy a very special form of unitary operation, namely phase shift, is assumed for encoding the information in the state of the probe. In fact, it has been shown that any disturbance to the considered unitary operation diminishes the precision De Pasquale et al. (2013). All these challenges may prevent the original interferometric sensing procedure to scale up for large system sizes or restrict it to special forms of interactions.
@@
An alternative method to interferometric quantum sensing has emerged in many-body systems. There are several distinct features between the two methods for sensing external parameters: (i) while the interferometric sensing requires entangled GHZ-type state preparation, in many-body sensors entanglement is either inherently present in the spectrum of the system or generated freely during the dynamics; (ii) while interaction between particles is destructive in interferometric sensing, it plays a crucial role in many-body sensors; (iii) unlike GHZ-type states, many-body systems are more resistive against decoherence and particle loss; and (iv) in contrast to the interferometric sensors, the optimal measurement may depend on the unknown parameter which is supposed to be measured. While features (i)-(iii) show the benefits of many-body sensors over their interferometric counterparts, the feature (iv) shows that taking advantage of many-body systems for sensing is not free of challenge. These pros and cons motivate further investigations of many-body systems for developing a new generation of sensors. In general, there are two different approaches to exploit many-body systems for metrology purposes with quantum enhanced sensitivity. The first approach, which typically requires equilibrium states, such as the ground state, exploits quantum criticality for achieving quantum enhanced sensitivity. In the second approach, however, non-equilibrium dynamics of a many-body system is exploited to accumulate information about the parameters of interest. Further divisions can be considered for non-equilibrium probes as they might be used at non-equilibrium steady states or transient dynamics. All these methods for achieving quantum enhanced sensitivity are summarized in Fig. 1.
@@
Figure 1: Schematic for achieving quantum enhanced sensitivity. The main advantage of quantum sensors over their classical counterparts is manifested through quantum enhanced sensitivity. Various strategies have been discovered to achieve such enhancement. Originally, interferometric quantum sensors were proposed and realized for measuring a phase shift with quantum enhanced precision. Later, the capacity of quantum many-body systems for achieving quantum enhanced sensitivity was identified which is the subject of this review. Many-body systems have been used for quantum sensing in both equilibrium (e.g. criticality in the ground state) and non-equilibrium modes. In the non-equilibrium case, we can further divide the sensors into non-equilibrium steady states and transient dynamics.
@@
Quantum criticality has already been identified as a resource for sensing in both equilibrium and non-equilibrium many-body systems. Originally, quantum phase transition in the transverse Ising model has been explored for metrology purposes Zanardi and Paunković (2006). It has been then extended to various types of criticalities including, second-order Zanardi and Paunković (2006); Zanardi et al. (2007); Gu et al. (2008); Zanardi et al. (2008); Invernizzi et al. (2008); Gu (2010); Gammelmark and Mølmer (2011); Skotiniotis et al. (2015); Rams et al. (2018); Wei (2019); Chu et al. (2021); Liu et al. (2021); Montenegro et al. (2021); Mirkhalaf et al. (2021); Di Candia et al. (2023); Mihailescu et al. (2024a), first-order Raghunandan et al. (2018); Heugel et al. (2019); Yang and Jacob (2019), topological Budich and Bergholtz (2020); Sarkar et al. (2022); Koch and Budich (2022); Yu et al. (2024a); Sarkar et al. (2024), non-Hermitian systems Chen et al. (2017); Yu et al. (2020); McDonald and Clerk (2020); Wiersig (2014); Hodaei et al. (2017); Liu et al. (2016); Langbein (2018); Chen et al. (2019); Wang et al. (2020); Zhang et al. (2019); Lau and Clerk (2018); Kononchuk et al. (2022); Wiersig (2020); Xiao et al. (2024); Yu et al. (2024b), and Stark phase transitions He et al. (2023); Yousefjani et al. (2023a, 2024a). Apart from these distinct types of criticalities at equilibrium, non-equilibrium quantum phase transitions can also demonstrate quantum enhanced sensitivity. This includes dissipative Fernández-Lorenzo and Porras (2017); Baumann et al. (2010); Baden et al. (2014); Klinder et al. (2015); Rodriguez et al. (2017); Fitzpatrick et al. (2017); Fink et al. (2017); Ilias et al. (2022, 2024); Alipour et al. (2014), Floquet Mishra and Bayat (2021, 2022) and time crystal Montenegro et al. (2023); Cabot et al. (2024); Lyu et al. (2020); Iemini et al. (2024a); Yousefjani et al. (2024b) phase transitions. Each type of quantum criticality comes with its own features and characteristics. For instance, second-order quantum phase transitions take place in the ground state of a many-body system and are accompanied by spontaneous symmetry breaking and long-range correlations which are described by Landau-Ginzburg formalism Sachdev (1999). On the other hand, topological phase transitions neither show symmetry breaking nor follow the Landau-Ginzburg theorem Xu (2012) and Stark localization takes place across the entire spectrum and not just the ground state van Nieuwenburg et al. (2019); Schulz et al. (2019). While different types of quantum criticalities have their own characteristics, one feature is common among all of them for showing quantum enhanced sensitivity which is energy/quasi-energy gap closing. This is an interesting observation which guides us in searching for finding new potential quantum sensors with the capability of achieving quantum enhanced precision. Interestingly, the realization of criticality-based quantum sensors are becoming viable in various physical platforms, including nuclear magnetic resonance systems Liu et al. (2021) and Rydberg atoms Ding et al. (2022).

### v3

@@ Abstract
Quantum systems, fabricated across various spatial scales from nano to micrometers, are very delicate and naturally sensitive to the variations of their environment. These features make them excellent candidates for serving as sensors with wide range of applications. Indeed, the exceptional precision of quantum sensors arises from their compact size and inherent sensitivity, enabling measurements with unprecedented accuracy within highly localized regions. A key advantage of quantum sensors lies in their resource efficiency, as their achievable precision can scale super-linearly with respect to resources, such as system size, in contrast to the linear scaling characteristic of classical sensors. This phenomenon, commonly referred to as quantum-enhanced sensitivity, fundamentally depends on exploiting uniquely quantum mechanical features, including superposition, entanglement, and squeezing. Originally, quantum sensing was formulated for particles prepared in a special form of entangled states. Yet, certain realization of these probes may be susceptible to decoherence and interaction between particles may also be detrimental to their performance. An alternative framework for quantum sensing has been developed through exploiting quantum many-body systems, where the interaction between particles plays a crucial role. In this review, we investigate different aspects of the latter approach for quantum metrology and sensing. Many-body probes have been used for sensing purposes in both equilibrium and non-equilibrium scenarios. Quantum criticality, as a well-studied subject in many-body physics, has been identified as a resource for achieving quantum-enhanced sensitivity in both of these scenarios. In equilibrium, various types of criticalities, such as first order, second order, topological, and localization phase transitions have been exploited for sensing purposes. In non-equilibrium scenarios, quantum-enhanced sensitivity has been discovered for Floquet, dissipative, and time crystal phase transitions. While each type of these criticalities, either in equilibrium or non-equilibrium scenarios, has its own characteristics, the presence of one feature is crucial for achieving quantum-enhanced sensitivity and that is energy/quasi-energy gap closing. In non-equilibrium quantum sensing, time becomes another parameter which can affect the sensitivity of the probe. Typically, the sensitivity enhances as the probe evolves in time. In this review, we provide an overview on recent progresses on different aspects of quantum metrology and sensing with many-body systems.
@@ I Introduction
The rapidly emerging quantum technologies are expected to revolutionize our lives in the coming decades Dowling and Milburn (2003). In a broad sense, quantum technologies are developed in three directions, namely quantum simulation and computation Lloyd (1996); Georgescu et al. (2014), quantum communication Zeilinger (1999); Gisin and Thew (2007), and quantum sensing Degen et al. (2017). Thanks to advancements in fabricating ultra-precise quantum devices, the progress in all these three aspects have significantly been accelerated. Quantum sensing Degen et al. (2017) is one of the pillars of quantum technologies which allow for developing a new generation of sensors for detecting gravitational, magnetic and electric fields as well as estimating physical quantities with unprecedented precision, well beyond the capability of conventional classical sensors. The application prospect of quantum sensors is very wide and includes mining Crawford et al. (2021), environmental monitoring  Lanzagorta et al. (2015), gravimetry Qvarfort et al. (2018, 2021); Armata et al. (2017); Montenegro (2025); Szigeti et al. (2020); Kritsotakis et al. (2018); Gietka et al. (2019), biological imaging Wu et al. (2016a); Taylor and Bowen (2016); Aslam et al. (2023), space exploration Kaltenbaek et al. (2021), radar and lidar technology Maccone and Ren (2020); Karsa et al. (2024); Casariego et al. (2023); Reichert et al. (2024), searching for fundamental particles Brady et al. (2022, 2023); Shi and Zhuang (2023); Backes et al. (2021); Kang et al. (2024); Wang et al. (2022a); Jiang et al. (2021a) — the list goes on. There are several reasons that make quantum sensors superior over their classical counterparts. First, the even smaller sizes of quantum sensors, which can be fabricated at the atomic scales, allows for measuring environmental parameters within an extremely localized spatial extensions, such as biological cells Schirhagl et al. (2014a); Zhang et al. (2021). Second, the natural delicacy of quantum superposition allows quantum sensors to detect very weak signals Hong et al. (2024); Jiang et al. (2021b). Third, quantum features, such as superposition, entanglement and squeezing, can be harnessed by quantum probes to achieve quantum-enhanced precision outperforming their classical counterparts using the same amount of resources Giovannetti et al. (2004, 2011, 2006). Fourth, quantum sensors have been developed on a wide variety of physical platforms, showing their flexibility in fabrication, control and application. Such platforms includes ultra-cold atoms Kasevich and Chu (1992); Peters et al. (1999); Bongs et al. (2019); El-Neaj et al. (2020); Stray et al. (2022); Peters et al. (2001); Fixler et al. (2007); Tino (2021); Bloom et al. (2014); Hinkley et al. (2013); Takamoto et al. (2005), ion traps Leibfried et al. (2004); Maiwald et al. (2009); Biercuk et al. (2010); Sawyer et al. (2012); Brownnutt et al. (2015); Baumgart et al. (2016, 2016); Huntemann et al. (2016), atomic vapors Budker and Romalis (2007); Kominis et al. (2003); Dang et al. (2010); Balabas et al. (2010); Shah et al. (2007); Fernholz et al. (2008); Wasilewski et al. (2010); Xia et al. (2006), nuclear magnetic resonance systems Kitching et al. (2011); Fang and Qin (2012); Jiang et al. (2021a); Wang et al. (2022a); Wu et al. (2022); Jiang et al. (2024), solid state defects in diamond Taylor et al. (2008); Clevenson et al. (2015); Jensen et al. (2014); Le Sage et al. (2012); Chernobrod and Berman (2005); Balasubramanian et al. (2008); Dolde et al. (2011); Zhao et al. (2012); Barson et al. (2017); Wu et al. (2021a); Vetter et al. (2022); Zhou et al. (2014); Xie et al. (2020); Schirhagl et al. (2014b); Patel et al. (2020), superconducting circuits Bal et al. (2012); Danilin et al. (2018); Wang et al. (2019a); Yu et al. (2025), photonic setups Holland and Burnett (1993); Mitchell et al. (2004); Pezzé et al. (2007); Higgins et al. (2007); Nagata et al. (2007); Ono et al. (2013); Xiao et al. (2024), and optomechanical devices Hu et al. (2024); Gavartin et al. (2012); Liu et al. (2022); Westerveld et al. (2021); Sansa et al. (2020); Li and Barker (2018); Fardian-Melamed et al. (2025); Fogliano et al. (2021); Pikovski et al. (2012); Chowdhury et al. (2023); Gosling et al. (2024); Forstner et al. (2012); McCormick et al. (2019); Aspelmeyer et al. (2014); Moser et al. (2013); Guzmán Cervantes et al. (2014); Krause et al. (2012); Chaste et al. (2012); Forstner et al. (2014); Degen et al. (2009); Rugar et al. (2004); Li et al. (2021); Liu et al. (2021a); Xia et al. (2023), the list goes on.
@@
In order to quantify the precision of a sensor for estimating an unknown parameter $\theta$, one needs to specify a proper figure of merit. In case of maximal-likelihood estimation theory, the imprecision of the estimation can be quantified by the variance of an unbiased max-likelihood estimator. This variance is bounded through Cramér-Rao inequality Cramer (1946); Rao (1992) by $1/MI(\theta)$, where $M$ is the number of samples and $I(\theta)$ is a quantity called Fisher information, which will be defined and discussed later. Dependence on $M$ is a direct consequence of central limit theorem, therefore the most important quantity which bounds the precision of a sensor is the Fisher information. In fact, in estimation theory, Fisher information is the conventional figure of merit for evaluating the performance of a sensor. Every sensing procedure exploits some resources such as time, particle number, system size, etc. The resource efficiency of a given sensor is determined by the scaling of Fisher information with respect to those resources. In other words, such scaling shows how the precision is enhanced by using more resources. In the absence of quantum features, the Fisher information at best scales linearly with the resource, which is known as standard quantum limit (or classical shot-noise limit). However, exploiting quantum features such as quantum coherence, entanglement, and squeezing may result in super-linear scaling of the Fisher information with respect to their resource. This is known as quantum-enhanced sensitivity and is one of the key properties that makes quantum sensors superior over their classical counterparts.
@@
Seminal ideas in quantum metrology can be traced back to the 1980s, when debates arose over whether quantum-mechanical radiation-pressure fluctuations disturb position measurements in an interferometry setup Caves (1980), as well as quantum mechanical noise in linear amplifiers Caves (1982) and interferometry Caves (1981). For further insights, see the theoretical review by Caves et al. Caves et al. (1980), along with an experimental progress review by Bocko et al. Bocko and Onofrio (1996). Moreover, in 1994, Samuel L. Braunstein and Carlton M. Caves Braunstein and Caves (1994) formulated a more general uncertainty principle based on a Riemannian metric on the space of quantum-mechanical density operators. Later, quantum-enhanced sensitivity was proposed in an interferometric setup by V. Giovannetti, S. Lloyd, and L. Maccone Giovannetti et al. (2004). In this paper, the authors show that a special form of entangled states, such as Greenberger–Horne–Zeilinger (GHZ) states, can enhance the precision of detecting a phase shift quadratically with respect to the probe size. Optical setups have since been used to experimentally confirm this effect quite extensively (Mitchell et al., 2004; Pezzé et al., 2007; Ono et al., 2013). Even for matter-based platforms, the effect was tested experimentally in ion-trap systems Leibfried et al. (2004) and then extended to various physical platforms like superconducting qubits Wang et al. (2019a) and nitrogen vacancy centers Bonato et al. (2016). The interferometry-based quantum sensing has several advantages, including: (i) the quantum-enhanced precision can be achieved for all range of parameters; and (ii) the measurement which results in quantum-enhanced precision is fixed and does not depend on the unknown parameter. Nonetheless, this approach also has inherent limitations, as it relies on GHZ-type entangled states, which are difficult to create and, in general, might be susceptible to decoherence and particle loss. In addition, in this strategy a very special form of unitary operation, namely phase shift, is assumed for encoding the information in the state of the probe. In fact, it has been shown that any disturbance to the considered unitary operation diminishes the precision De Pasquale et al. (2013). All these challenges may prevent the original interferometric sensing procedure to scale up for large system sizes or restrict it to special forms of interactions.
@@
An alternative method to interferometric quantum sensing has emerged in many-body systems. There are several distinct features between the two methods for sensing external parameters: (i) while the interferometric sensing requires entangled GHZ-type state preparation, in many-body sensors entanglement is either inherently present in the spectrum of the system or generated freely during the dynamics; (ii) while interaction between particles is usually destructive in interferometric sensing De Pasquale et al. (2013), it plays a crucial role in many-body sensors; (iii) unlike GHZ-type states, many-body systems are more resistive against decoherence and particle loss; and (iv) in contrast to the interferometric sensors, the optimal measurement may depend on the unknown parameter which is supposed to be measured. While features (i)-(iii) show the benefits of many-body sensors over their interferometric counterparts, the feature (iv) shows that taking advantage of many-body systems for sensing is not free of challenge. These pros and cons motivate further investigations of many-body systems for developing a new generation of sensors. See, for instance, Ref. Baak and Fischer (2022) for an analytically solvable many-body problem in Hamiltonian parameter estimation, both for classical and quantum metrology. In general, there are two different approaches to exploit many-body systems for metrology purposes with quantum-enhanced sensitivity. The first approach, which typically requires equilibrium states, such as the ground state, exploits quantum criticality for achieving quantum-enhanced sensitivity. In the second approach, however, non-equilibrium dynamics of a many-body system is exploited to accumulate information about the parameters of interest. Further divisions can be considered for non-equilibrium probes as they might be used at non-equilibrium steady states or transient dynamics. All these methods for achieving quantum-enhanced sensitivity are summarized in Fig. 1. The advantages and limitations of each of these methods are comprehensively summarized in Table 1. Additionally, the frequently used notations in this review article are summarized in Table 2 for convenience.
@@
Figure 1: Schematic for achieving quantum-enhanced sensing. The main advantage of quantum sensors over their classical counterparts is manifested through quantum-enhanced sensitivity. Various strategies have been discovered to achieve such enhancement. Originally, interferometric quantum sensors were proposed and realized for measuring a phase shift with quantum-enhanced precision. Later, the capacity of quantum many-body systems for achieving quantum-enhanced sensitivity was identified which is the subject of this review. Many-body systems have been used for quantum sensing in both equilibrium (e.g. criticality in the ground state) and non-equilibrium modes. In the non-equilibrium case, we can further divide the sensors into non-equilibrium steady states and transient dynamics.
@@
Quantum criticality has already been identified as a resource for sensing in both equilibrium and non-equilibrium many-body systems. Originally, quantum phase transition in the transverse Ising model has been explored for metrology purposes Zanardi and Paunković (2006). It has been then extended to various types of criticalities including, second-order Zanardi and Paunković (2006); Zanardi et al. (2007); Gu et al. (2008); Zanardi et al. (2008); Invernizzi et al. (2008); Gu (2010); Gammelmark and Mølmer (2011); Skotiniotis et al. (2015); Rams et al. (2018); Wei (2019); Chu et al. (2021); Liu et al. (2021b); Montenegro et al. (2021); Mirkhalaf et al. (2021); Di Candia et al. (2023); Mihailescu et al. (2024a), first-order Raghunandan et al. (2018); Mirkhalaf et al. (2020); Yang and Jacob (2019); Sarkar et al. (2025), topological Budich and Bergholtz (2020); Sarkar et al. (2022); Koch and Budich (2022); Yu et al. (2024a); Sarkar et al. (2024), and Stark phase transitions He et al. (2023); Yousefjani et al. (2023a, 2025a); Sarkar and Bayat (2025). Apart from these distinct types of criticalities at equilibrium, non-equilibrium quantum phase transitions can also demonstrate quantum-enhanced sensitivity. This includes dissipative Fernández-Lorenzo and Porras (2017); Baumann et al. (2010); Baden et al. (2014); Klinder et al. (2015); Rodriguez et al. (2017); Fitzpatrick et al. (2017); Fink et al. (2017); Ilias et al. (2022, 2024); Alipour et al. (2014a), Floquet Mishra and Bayat (2021, 2022) and time crystal Montenegro et al. (2023); Cabot et al. (2024); Lyu et al. (2020); Iemini et al. (2024a); Yousefjani et al. (2025b) phase transitions. Each type of quantum criticality comes with its own features and characteristics. For instance, second-order quantum phase transitions take place in the ground state of a many-body system and are accompanied by spontaneous symmetry breaking and long-range correlations which are described by Landau-Ginzburg formalism Sachdev (1999). On the other hand, topological phase transitions neither show symmetry breaking nor follow the Landau-Ginzburg theorem Xu (2012) and Stark localization takes place across the entire spectrum and not just the ground state van Nieuwenburg et al. (2019); Schulz et al. (2019). While different types of quantum criticalities have their own characteristics, one feature is common among all of them for showing quantum-enhanced sensitivity which is energy/quasi-energy gap closing. This is an interesting observation which guides us in searching for finding new potential quantum sensors with the capability of achieving quantum-enhanced precision. Interestingly, the realization of criticality-based quantum sensors are becoming viable in various physical platforms, including nuclear magnetic resonance systems Liu et al. (2021b),Rydberg atoms Ding et al. (2022) and superconducting devices Yu et al. (2025).
Sensing Strategy Advantages Limitations
Interferometric Quantum Sensors
External parameter is
induced by phase shift
$\hat{U}(\theta)=e^{i\theta\hat{G}}$
1. Heisenberg scaling possible
2. Universal optimal measurement
1. Difficult preparation
2. Potentially sensitive to decoherence
and particle loss
3. Works only for phase shift operations
Many-Body Quantum Sensors at Equilibrium
Ground or Thermal
State based for Hamiltonians
of the form $\hat{H}{=}\hat{H}_{0}{+}\hat{H}_{1}(\theta)$
1. Quantum-enhanced sensitivity achievable
by gap-closing phase transitions
2. Quantum-enhanced sensitivity possible
3. Robust against imperfections
1. Fine tuning is needed (local sensing)
2. Optimal measurement parameter-value
dependent
3. Preparation might be resource consuming
Non-Equilibrium Many-Body Quantum Sensors
Non-Equilibrium Steady
State based
1. Initialization does not matter
2. Quantum-enhanced sensitivity possible
3. Works even with partial accessibility
4. Dissipation can also be harnessed
1. Fine tuning is needed (local sensing)
2. Optimal measurement parameter-value
dependent
3. Reaching steady state is time consuming
Transient State based
1. Easy initialization
2. Quantum-enhanced sensitivity possible
3. Time control for better precision
4. May operate across the entire phase
(no fine tuning)
5. Sequential measurements are adaptable
1. Decoherence is a limiting factor
2. Optimal measurement parameter-value
dependent
Notation Meaning
$I(\theta)$ Classical Fisher information (or just Fisher information) with respect to unknown parameter $\theta$.
$I_{Q}(\theta)$ Quantum Fisher information with respect to unknown parameter $\theta$.
$\tilde{\Theta}$ Statistical estimator corresponding to unknown parameter $\theta$.
$\partial_{\theta}$ Partial derivative with respect to unknown parameter $\theta$.
$F(\rho_{1},\rho_{2})$ Quantum fidelity between two density matrices $\rho_{1}$ and $\rho_{2}$, given by $F(\rho_{1},\rho_{2})=\sqrt{\text{Tr}\sqrt{\sqrt{\rho_{1}}\rho_{2}\sqrt{\rho_{1}}}}$.
$\mathrm{Var}[\cdot]$ Variance of a random variable.
$\hat{O}$ Any quantum mechanical operator.
$\bm{\theta}$ Vector of multiple parameters.
$N_{\rm ex}$ Number of particles (or excitations) in a many-body system.
$L$ and $N$ Depending on the context, both have been used for the size of a many-body system.
$\hat{\sigma}_{\cdot}$ $2\times 2$ Pauli matrices.
$n_{\mathrm{seq}}$ Number of consecutive steps in sequential measurement sensing protocols.
$T$ Temperature.
$B$ and $h$ Magnetic fields.
Table 2: Notation used throughout the review, unless specified otherwise.